OpenStudy (anonymous):

what is the slope of the line that passes through 3,2 and 5,6

OpenStudy (anonymous):

(6-2)/(5-3)

OpenStudy (anonymous):

4/2

OpenStudy (anonymous):

2

OpenStudy (anonymous):

do you know where to go from here? y=2x+b

OpenStudy (anonymous):

(2)=2(3) + b 2=6+b b=-4

OpenStudy (anonymous):

so the equation for a line that passes through those points is y=2x-4

OpenStudy (anonymous):

thank u